You have to mix sealant that has a ratio of 16:100, catalyst to base. How much catalyst and how much base must you use to make 638 grams of sealant?
Answer by RAY100(1637)

c/b=16/100,,,,or c=.16 b
.
c+b= 638
.
(.16b) +b = 638
.
1.16b = 638
.
b=550,,,,,,,,,,,,,,c=.16(550 = 88
.
.
check 88/550 =.16,,,,,,,ok
.
88+550 = 638 ,,,,,ok